OpenStudy (anonymous):

write te slope-intercept form of the line that passes through (1,-7) and (5,-19)

OpenStudy (anonymous):

slope intercept form: \[y = mx+b \]first find the slope of the line using slope formula of \[\frac{y_2-y_1}{x_2-x_1} \]then put it in point slope form. which is \[y-y_1 = m(x-x_1) \]then you can turn that into\[ y=mx+b \] form